# A Survey on RAG Meeting LLMs: Towards Retrieval-Augmented Large Language Models

### v2

@@ 1. Introduction
As one of the most fundamental data mining techniques, retrieval aims to understand the input query and extract relevant information from external data sources (Kobayashi and Takeda, 2000; Singhal et al., 2001). It has found extensive application in various fields (Buttcher et al., 2016; Yin et al., 2016; O’Hare et al., 2016), such as search, question answering, and recommender systems. For instance, search engines (e.g., Google, Bing, and Baidu) are the most successful applications of retrieval in the industry; they can filter and retrieve the most relevant web pages or documents that can match a user’s query (Croft et al., 2010; Yin et al., 2016), enabling users to find the desired information effectively. Meanwhile, retrieval models, through effective data maintenance in external databases, can provide faithful and timely external knowledge, thereby serving vital functions in various knowledge-intensive tasks. Due to their powerful capacities, retrieval techniques have been successfully incorporated into advanced generative models in the era of AI-Generated Content (AIGC) (Li et al., 2023a; Wu et al., 2024; Sheynin et al., 2023; Zhang et al., 2023a). Notably, the integration of retrieval models with language models has given rise to Retrieval-Augmented Generation (RAG) (Lewis et al., 2020c), which has emerged as one of the most representative techniques in the field of generative AI, aiming to enhance the generation quality of text content with retrieved information (Li et al., 2023a; Lewis et al., 2020c; Borgeaud et al., 2022).
@@
Recent years have witnessed the rapid development of pre-trained foundation models, particularly Large Language Models (LLMs), which have demonstrated impressive performance across various tasks (Chowdhery et al., 2023; Achiam et al., 2023), including recommender systems (Zhao et al., 2024a), molecule discovery (Li et al., 2023a), and report generation (Ding et al., 2024). The great success of LLMs can be technically attributed to the advanced architectures with billion-level parameters pre-training on a huge amount of training corpus from various sources. These technical improvements have given rise to the remarkable emergence capabilities of LLMs (Zhao et al., 2023b, 2024a), particularly in language understanding and generation, in-context learning, and others. For instance, GPT-FAR introduces detailed prompts to teach GPT-4 to perform image tagging, statistical analysis, and text analysis for multi-modal fashion report generation (Ding et al., 2024). LLMs also achieve promising performance in recommender systems by understanding users’ preferences towards items (Wang et al., 2024a; Zhao et al., 2024a). Despite the success, LLMs still suffer from intrinsic limitations (Zhao et al., 2024a, 2023b), such as the lack of domain-specific knowledge, the problem of “hallucination”, and the substantial computational resources required for updating the models. These problems are particularly notable in domain-specific fields like medicine and law. For instance, a recent study has demonstrated that legal hallucinations are pervasive and disturbing, with hallucination rates ranging from 69% to 88% in responses to specific legal queries for state-of-the-art LLMs (Dahl et al., 2024). Moreover, the challenges of tackling the hallucination problem become even harder due to the substantial computational resources required for fine-tuning LLMs with domain-specific or the latest data. This, in turn, significantly hinders the widespread adoption of LLMs in various real-world applications.
@@ 2.1. Large Language Models (LLMs)
Figure 2. Representing RAG and RA-LLMs methods organized by their main design focus, proposed time and impact (shown by citation). Note that the year shown along with the model name matches with the reference while the arrangement of the these models is based on their first time being published.
@@ 2.2.2. In-Context Learning (ICL)
To overcome the limitations of vanilla prompt learning, recent efforts (Liu et al., 2022a; Kim et al., 2022; Zhang et al., 2023d) have developed in-context learning (ICL). ICL is a specific method of prompt learning that gives the model a few demonstrations of tasks within the prompt. This paradigm allows pre-trained LLMs to understand the pattern provided by the demonstrations to solve novel tasks without the need for fine-tuning. For example, by carefully selecting a few demonstrations, GPT-3 (Brown et al., 2020) has shown the capability to perform few-shot tasks (Liu et al., 2022a). This success indicates that LLMs have a remarkable ability to rapidly adapt to new tasks based on task-specific knowledge.
Despite its effectiveness, ICL usually relies heavily on the quality of the provided demonstrations (Liu et al., 2022b; Su et al., 2023), which may lead to the generation of sub-optimal outputs. Even worse, ICL may not have enough necessary information or prior knowledge to guide the LLMs in generating accurate responses. To address the aforementioned limitations of ICL, more recent studies introduce Retrieval-Augmented Generation (RAG) technologies (Lewis et al., 2020c; Ram et al., 2023; Shi et al., 2023). By integrating retrieval with generation, RAG models provide a promising direction for enhancing the performance and adaptability of LLMs across various tasks.
@@ 3.1.3. Pre-retrieval and Post-retrieval Enhancement
Table 1. Basic information and main technical designs of high-impact RAG and RA-LLM models.1
@@ 3.3.2. Output-Layer Integration
Another kind of augmentation is post-hoc, i.e., output-layer integration, which joints retrieval and generation results. For example, kNN-LM (Khandelwal et al., 2020) interpolates two next-token distributions in prediction: one induced by the LM and the other induced by the nearest neighbors from the retrieval corpus. Output-layer linear integration (Grave et al., 2017; Zhong et al., 2022) is flexible to apply since it can be plugged into most generation models without additional training. However, the simplicity of output-layer integration also limits the model’s ability to reason about the retrieved text. To tackle this limitation, Yogatama et al. (2021) propose to add an extra gating network to post-process the retrieved data and achieve comparatively better performance. For LLMs, output-layer integration is as reasonable and adaptive as input-layer integration. REFEED (Yu et al., 2023c) proposes an answer refining mechanism that applies an LLM to evaluate the retrieved information and adjust the initial answer accordingly to enhance the accuracy of the response. Similarly, Zhang et al. (2023b) propose the COMBO framework, which matches LLM-generated passages with retrieved counterparts into compatible pairs based on pre-trained discriminators. The passage pairs are then handled by a Fusion-in-Decoder-based (Izacard and Grave, 2021b) to derive a final answer.
@@ 4.3.1. Retriever First
Applications NLP Applications QA Systems ChatBots Fact Verification Downstream Tasks Recommendations Software Engineering Domain-specific Applications AI for Science Finance RETRO (Borgeaud et al., 2022) Fusion-in-Decoder (Izacard and Grave, 2021b) REALM (Guu et al., 2020), etc. Ghazvininejad et al. (2018) KDBTS (Chen et al., 2020) Komeili et al. (2022) Wang et al. (2023d), etc. RAG (Lewis et al., 2020c) Atlas (Izacard et al., 2023) Self-RAG (Asai et al., 2023b), etc. Yepes et al. 2024 Di Palma (2023) CoRAL (Wu et al., 2024) RevCore (Lu et al., 2021), etc. Clinfo. ai (Lozano et al., 2023) RetMol (Wang et al., 2023b) MoleculeSTM (Liu et al., 2023) PMINet (Yang et al., 2023a) BioBridge (Wang et al., 2023e) RSA (Ma et al., 2023b) Graphvf (Sun et al., 2023), etc. Docprompting (Zhou et al., 2022) Atlas (Nashid et al., 2023) Dater (Ye et al., 2023a) SheetCopilo (Li et al., 2024b) Xricl (Shi et al., 2022) Synchromesh (Poesia et al., 2022), etc. Zhang et al. (2023c) AlphaFin (Li et al., 2024a) ChatDOC (Lin, 2024) Yepes et al. (2024), etc.
@@ 4.4. Joint Training
Joint training methods (Zhong et al., 2022; Kang et al., 2023; Li et al., 2023b; Xu et al., 2023c; Hu et al., 2023; Cheng et al., 2024) employ the end-to-end paradigm to optimize the retriever and generator simultaneously. Instead of training each module sequentially, joint training methods effectively enhance the retriever’s ability to locate external knowledge for generation and the generator’s capacity to effectively leverage the retrieved information. For instance, RAG (Lewis et al., 2020c) minimizes the negative loglikelihood to jointly train the retriever and generator. REALM (Guu et al., 2020) adopts a similar training paradigm to that of RAG (Lewis et al., 2020c), and Maximum Inner Product Search (MIPS) (Ram and Gray, 2012; Chen et al., 2019; Shen et al., 2015; Ding et al., 2020) technique is used to locate the most relevant documents. To employ MIPS, all external documents are embedded first and a search index is produced for each embedding. An asynchronous index updating strategy (Guu et al., 2020; Izacard et al., 2023; Siriwardhana et al., 2023; Huang et al., 2023) is proposed to refresh the index every several hundred training steps to avoid time consumption of re-indexing all documents.
@@ 5.1.1. QA Systems
QA Systems aim to provide precise answers to user’s queries. However, even when trained on extensive data, these systems may lack the latest information or specific domain knowledge that is not included in their training data (Izacard and Grave, 2021b; Liu et al., 2022c). To address this limitation, the integration of RA-LLMs has played a crucial role in advancing the capabilities of QA systems by enhancing their ability to retrieve and synthesize relevant information (Borgeaud et al., 2022; Izacard and Grave, 2021b). Specifically, RA-LLMs can provide coherent and contextually relevant answers by leveraging their retrieval component to access a vast knowledge base. For example, REALM (Guu et al., 2020) integrates a knowledge retriever that can retrieve information from a large corpus during pre-training, fine-tuning, and inference. This approach allows REALM to effectively retrieve from a vast knowledge corpus, thereby improving the accuracy of its responses. Similarly, Fusion-in-Decoder (Izacard and Grave, 2021b) retrieves passages from support documents and then fuses them with questions to generate the answer, achieving higher accuracy. In addition,  Borgeaud et al. (2022) indicate that the quality of the answers may rely more on the output of retrieval.
@@ 5.3.2. Finance
In the highly data-driven and information-intensive field of finance, RA-LLMs have proved to be a significant technology for enhancing decision-making (Zhang et al., 2023c; Yepes et al., 2024; Li et al., 2024a). For example, Zhang et al. (2023c) retrieve financial information from external sources, such as news platforms (e.g., Bloomberg and Reuters) and social media platforms (e.g., Twitter, Reddit), to combine with the original query to enhance the precision of financial sentiment analysis. In addition, financial QA is another primary task of financial analysis, which usually extracts relevant knowledge from financial documents. As professional documents are usually stored in PDF format, Lin (2024) introduces a PDF parser combined with RA-LLMs to retrieve knowledge from financial reports. On the other hand, Yepes et al. (2024) propose a document chunking method based on structure instead of chunking based on paragraphs, further improving the quality of RA-LLMs outputs.
@@ 6. Future Challenges and Opportunities
Trustworthy RA-LLMs. The essential objective of developing RAG-empowered LLMs is to enhance the capability of the language models, thereby benefiting users and society by alleviating redundant and meaningless labor, increasing conveniences, and spurring social progress. However, recent research indicates that RA-LLMs can be maliciously and unintentionally manipulated to make unreliable decisions and harm humans (Deng et al., 2024; Zou et al., 2024), which may have serious consequences in safety-critical scenarios (Liu et al., 2021; Fan et al., 2022b, 2021; Chen et al., 2023b, 2022). In addition, private retrieval database has a risk of leakage, raising concerns regarding the privacy of RA-LLMs (Zeng et al., 2024). Therefore, developing trustworthy RA-LLMs is of paramount importance as it can significantly mitigate the potential negative impacts of LLMs technology and provide people with powerful AI models that can be fully trusted. To be specific, the ideal trustworthiness in RA-LLMs systems should possess the following characteristics: 1) robustness, 2) fairness, 3) explainability, and 4) privacy. For example, robustness means a trustworthy RA-LLMs system should be robust against malicious or inadvertent perturbations introduced by attackers. Fairness indicates a trustworthy RA-LLMs system ought to avoid discrimination during the decision-making process. Explainability requires a complete understanding of the intrinsic workings of RA-LLMs systems, i.e., the predictions of RA-LLMs systems are explainable and transparent. Privacy entails safeguarding the safety of this private information housed within the datastore when establishing trustworthy RA-LLMs systems.

### v3

@@ 1. Introduction
As one of the most fundamental data mining techniques, retrieval aims to understand the input query and extract relevant information from external data sources (Kobayashi and Takeda, 2000; Singhal et al., 2001; Deng et al., 2024a; Ding et al., 2020b). It has found extensive application in various fields (Buttcher et al., 2016; Yin et al., 2016; O’Hare et al., 2016; Ding et al., 2023), such as search, question answering, and recommender systems. For instance, search engines (e.g., Google, Bing, and Baidu) are the most successful applications of retrieval in the industry; they can filter and retrieve the most relevant web pages or documents that can match a user’s query (Croft et al., 2010; Yin et al., 2016), enabling users to find the desired information effectively. Meanwhile, retrieval models, through effective data maintenance in external databases, can provide faithful and timely external knowledge, thereby serving vital functions in various knowledge-intensive tasks. Due to their powerful capacities, retrieval techniques have been successfully incorporated into advanced generative models in the era of AI-Generated Content (AIGC) (Li et al., 2023a; Wu et al., 2024; Sheynin et al., 2023). Notably, the integration of retrieval models with language models has given rise to Retrieval-Augmented Generation (RAG) (Lewis et al., 2020c), which has emerged as one of the most representative techniques in the field of generative AI, aiming to enhance the quality of the generated text content with retrieved information (Li et al., 2023a; Lewis et al., 2020c; Borgeaud et al., 2022).
@@
Recent years have witnessed the rapid development of pre-trained foundation models, particularly Large Language Models (LLMs), which have demonstrated impressive performance across various tasks (Chowdhery et al., 2023; Achiam et al., 2023), including recommender systems (Zhao et al., 2024a), molecule discovery (Li et al., 2023a), and report generation (Ding et al., 2024). Technically, the great success of LLMs can be technically attributed to the advanced architectures with billion-level parameters pre-training on a huge amount of training corpus from various sources. These technical improvements have given rise to the remarkable emergence capabilities of LLMs (Zhao et al., 2023b, 2024a), particularly in language understanding and generation, in-context learning, and others. For instance, GPT-FAR introduces detailed prompts to teach GPT-4 to perform image tagging, statistical analysis, and text analysis for multi-modal fashion report generation (Ding et al., 2024). LLMs also achieve promising performance in recommender systems by understanding users’ preferences towards items (Wang et al., 2024a; Zhao et al., 2024a). Despite the success, LLMs still suffer from intrinsic limitations (Zhao et al., 2024a, 2023b), such as the lack of domain-specific knowledge, the problem of “hallucination”, and the substantial computational resources required for updating the models. These problems are particularly notable in domain-specific fields like medicine and law. For instance, a recent study has demonstrated that legal hallucinations are pervasive and disturbing, with hallucination rates ranging from 69% to 88% in responses to specific legal queries for state-of-the-art LLMs (Dahl et al., 2024). Moreover, the challenges of tackling the hallucination problem become even harder due to the substantial computational resources required for fine-tuning LLMs with domain-specific or the latest data. This, in turn, significantly hinders the widespread adoption of LLMs in various real-world applications.
@@ 2.1. Large Language Models (LLMs)
Figure 2. Representing RAG and RA-LLMs methods organized by their main design focus, proposed time and impact (shown by citation). Note that the first author and year shown in the figure along with the model name can be used to locate corresponding reference.
@@ 2.2.2. In-Context Learning (ICL)
To overcome the limitations of vanilla prompt learning, recent efforts (Liu et al., 2022a; Kim et al., 2022; Zhang et al., 2023c) have developed in-context learning (ICL). ICL is a specific method of prompt learning that gives the model a few demonstrations of tasks within the prompt. This paradigm allows pre-trained LLMs to understand the pattern provided by the demonstrations to solve novel tasks without the need for fine-tuning. For example, by carefully selecting a few demonstrations, GPT-3 (Brown et al., 2020) has shown the capability to perform few-shot tasks (Liu et al., 2022a). This success indicates that LLMs have a remarkable ability to rapidly adapt to new tasks based on task-specific knowledge.
Despite its effectiveness, ICL usually relies heavily on the quality of the provided demonstrations (liu2022makes; Su et al., 2023), which may lead to the generation of sub-optimal outputs. Even worse, ICL may not have enough necessary information or prior knowledge to guide the LLMs in generating accurate responses. To address the aforementioned limitations of ICL, more recent studies introduce Retrieval-Augmented Generation (RAG) technologies (Lewis et al., 2020c; Ram et al., 2023; Shi et al., 2023). By integrating retrieval with generation, RAG models provide a promising direction for enhancing the performance and adaptability of LLMs across various tasks.
@@ 3.1.3. Pre-retrieval and Post-retrieval Enhancement
Table 1. Basic publication information and main technical designs of high-impact RAG and RA-LLM models.1
@@ 3.3.2. Output-Layer Integration
Another kind of augmentation is post-hoc, i.e., output-layer integration, which joints retrieval and generation results. For example, kNN-LM (Khandelwal et al., 2020) interpolates two next-token distributions in prediction: one induced by the LM and the other induced by the nearest neighbors from the retrieval corpus. Output-layer linear integration (Grave et al., 2017; Zhong et al., 2022) is flexible to apply since it can be plugged into most generation models without additional training. However, the simplicity of output-layer integration also limits the model’s ability to reason about the retrieved text. To tackle this limitation, Yogatama et al. (2021) propose to add an extra gating network to post-process the retrieved data and achieve comparatively better performance. For LLMs, output-layer integration is as reasonable and adaptive as input-layer integration. REFEED (Yu et al., 2023c) proposes an answer refining mechanism that applies an LLM to evaluate the retrieved information and adjust the initial answer accordingly to enhance the accuracy of the response. Similarly, Zhang et al. (2023a) propose the COMBO framework, which matches LLM-generated passages with retrieved counterparts into compatible pairs based on pre-trained discriminators. The passage pairs are then handled by a Fusion-in-Decoder-based (Izacard and Grave, 2021b) to derive a final answer.
@@ 4.3.1. Retriever First
Applications NLP Applications QA Systems ChatBots Fact Verification Downstream Tasks Recommendations Software Engineering Domain-specific Applications AI for Science Finance RETRO (Borgeaud et al., 2022) Fusion-in-Decoder (Izacard and Grave, 2021b) REALM (Guu et al., 2020), etc. Ghazvininejad et al. (2018) KDBTS (Chen et al., 2020) Komeili et al. (2022) Wang et al. (2023d), etc. RAG (Lewis et al., 2020c) Atlas (Izacard et al., 2023) Self-RAG (Asai et al., 2023b), etc. Yepes et al. 2024 Di Palma (2023) CoRAL (Wu et al., 2024) RevCore (Lu et al., 2021), etc. Clinfo. ai (Lozano et al., 2023) RetMol (Wang et al., 2023b) MoleculeSTM (Liu et al., 2023) PMINet (Yang et al., 2023a) BioBridge (Wang et al., 2023e) RSA (Ma et al., 2023b) Graphvf (Sun et al., 2023), etc. Docprompting (Zhou et al., 2022) Atlas (Nashid et al., 2023) Dater (Ye et al., 2023a) SheetCopilo (Li et al., 2024b) Xricl (Shi et al., 2022) Synchromesh (Poesia et al., 2022), etc. Zhang et al. (2023b) AlphaFin (Li et al., 2024a) ChatDOC (Lin, 2024) Yepes et al. (2024), etc.
@@ 4.4. Joint Training
Joint training methods (Zhong et al., 2022; Kang et al., 2023; Li et al., 2023b; Xu et al., 2023c; Hu et al., 2023; Cheng et al., 2024) employ the end-to-end paradigm to optimize the retriever and generator simultaneously. Instead of training each module sequentially, joint training methods effectively enhance the retriever’s ability to locate external knowledge for generation and the generator’s capacity to effectively leverage the retrieved information. For instance, RAG (Lewis et al., 2020c) minimizes the negative loglikelihood to jointly train the retriever and generator. REALM (Guu et al., 2020) adopts a similar training paradigm to that of RAG (Lewis et al., 2020c), and Maximum Inner Product Search (MIPS) (Ram and Gray, 2012; Chen et al., 2019; Shen et al., 2015; Ding et al., 2020a) technique is used to locate the most relevant documents. To employ MIPS, all external documents are embedded first and a search index is produced for each embedding. An asynchronous index updating strategy (Guu et al., 2020; Izacard et al., 2023; Siriwardhana et al., 2023; Huang et al., 2023) is proposed to refresh the index every several hundred training steps to avoid time consumption of re-indexing all documents.
@@ 5.1.1. QA Systems
QA Systems aim to provide precise answers to user’s queries. However, even when trained on extensive data, these systems may lack the latest information or specific domain knowledge that is not included in their training data (Izacard and Grave, 2021b; Liu et al., 2022b). To address this limitation, the integration of RA-LLMs has played a crucial role in advancing the capabilities of QA systems by enhancing their ability to retrieve and synthesize relevant information (Borgeaud et al., 2022; Izacard and Grave, 2021b). Specifically, RA-LLMs can provide coherent and contextually relevant answers by leveraging their retrieval component to access a vast knowledge base. For example, REALM (Guu et al., 2020) integrates a knowledge retriever that can retrieve information from a large corpus during pre-training, fine-tuning, and inference. This approach allows REALM to effectively retrieve from a vast knowledge corpus, thereby improving the accuracy of its responses. Similarly, Fusion-in-Decoder (Izacard and Grave, 2021b) retrieves passages from support documents and then fuses them with questions to generate the answer, achieving higher accuracy. In addition,  Borgeaud et al. (2022) indicate that the quality of the answers may rely more on the output of retrieval.
@@ 5.3.2. Finance
In the highly data-driven and information-intensive field of finance, RA-LLMs have proved to be a significant technology for enhancing decision-making (Zhang et al., 2023b; Yepes et al., 2024; Li et al., 2024a). For example, Zhang et al. (2023b) retrieve financial information from external sources, such as news platforms (e.g., Bloomberg and Reuters) and social media platforms (e.g., Twitter, Reddit), to combine with the original query to enhance the precision of financial sentiment analysis. In addition, financial QA is another primary task of financial analysis, which usually extracts relevant knowledge from financial documents. As professional documents are usually stored in PDF format, Lin (2024) introduces a PDF parser combined with RA-LLMs to retrieve knowledge from financial reports. On the other hand, Yepes et al. (2024) propose a document chunking method based on structure instead of chunking based on paragraphs, further improving the quality of RA-LLMs outputs.
@@ 6. Future Challenges and Opportunities
Trustworthy RA-LLMs. The essential objective of developing RAG-empowered LLMs is to enhance the capability of the language models, thereby benefiting users and society by alleviating redundant and meaningless labor, increasing conveniences, and spurring social progress. However, recent research indicates that RA-LLMs can be maliciously and unintentionally manipulated to make unreliable decisions and harm humans (Deng et al., 2024b; Zou et al., 2024), which may have serious consequences in safety-critical scenarios (Liu et al., 2021; Fan et al., 2022b, 2021; Chen et al., 2023b, 2022). In addition, private retrieval database has a risk of leakage, raising concerns regarding the privacy of RA-LLMs (Zeng et al., 2024). Therefore, developing trustworthy RA-LLMs is of paramount importance as it can significantly mitigate the potential negative impacts of LLMs technology and provide people with powerful AI models that can be fully trusted. To be specific, the ideal trustworthiness in RA-LLMs systems should possess the following characteristics: 1) robustness, 2) fairness, 3) explainability, and 4) privacy. For example, robustness means a trustworthy RA-LLMs system should be robust against malicious or inadvertent perturbations introduced by attackers. Fairness indicates a trustworthy RA-LLMs system ought to avoid discrimination during the decision-making process. Explainability requires a complete understanding of the intrinsic workings of RA-LLMs systems, i.e., the predictions of RA-LLMs systems are explainable and transparent. Privacy entails safeguarding the safety of this private information housed within the datastore when establishing trustworthy RA-LLMs systems.